# LLMs for Extremely Low-Resource Finno-Ugric Languages

### v1

@@ 1 Introduction
We present a case study in developing LLMs for extremely low-resource (XLR) languages, focusing on three Finno-Ugric (smugri [^1] Finno-Ugric translates to Estonian as soome-ugri, to Finnish as suomalais-ugrilaiset, to Võro as soomõ-ugri, and to Livonian as sūomõ-ugrõ, which is why we refer to it as smugri.) languages: Livonian, Võro and Komi. According to Joshi et al. (2020), Livonian is in the lowest category out of 6 (0 / The Left-Behinds) while Võro and Komi are in the second-lowest (1 / The Scrabing-Bys). Developing LLMs and other tools for these languages is thus a significant challenge, as well as a vital step for ensuring their digital survival and support.
@@ 3.1 Continued Pre-training
Stage 2 + parallel: making use of parallel translation data. To investigate the role of parallel translation examples in the pre-training data, we incorporate translation examples formatted into various templates, accounting for up to 1% of the Stage 2 character budget (159,712 sentence pairs). We use Unimax with N=1 to balance the budget between language pairs. For further details on the parallel data, we refer the reader to Appendix H.
@@ 3.2 Instruction-tuning
Supporting instructions. We utilize existing instruction-tuning datasets across multiple languages. For English, Russian, and Finnish, we use Aya Singh et al. (2024), and the highest-rated conversation paths of OASST-2 Köpf et al. (2023). Additionally, we sample 20,000 Estonian instructions from Alpaca-est Kuulmets et al. (2024). Following Kuulmets et al. (2024), we include 5,000 instructions from the FLAN-V2 Longpre et al. (2023) TULU mixture Wang et al. (2023) and 20,000 examples from Alpaca-GPT-4 Peng et al. (2023), to improve cross-lingual knowledge transfer from high-quality English instructions. We refer to this instruction mixture as SupInst (Supporting Instructions). Further details are listed in Appendix F
@@
To investigate a scenario where a translation model is unavailable, we explore handling translating Alpaca instructions to XLR languages by fine-tuning the based model for the translation task (discussed in §3.3) resulting in LLMTrAlpaca instructions. This is similar to self-translate-train Ri et al. (2024) and self-translate Etxaniz et al. (2024a) concepts with the difference that we add a fine-tuning step to obtain an LLM specialized for translation. Further instruction-translation details are discussed in Appendix G.
@@ 4.2.2 Constructing the Dataset
Model SIB-SMUGRI BELEBELE-SMUGRI FLORES-SMUGRI byte-PPL
@@
VRO LIV KPV VRO LIV KPV ET-VRO ET-LIV RU-KPV VRO LIV KPV
Llammas-base 78.4 ± 3.7 69.6 ± 4.1 64.0 ± 4.3 30.7 ± 4.1 28.4 ± 4.0 32.3 ± 4.2 11.5 ± 0.9 4.3 ± 0.5 1.7 ± 0.4 3.3548 12.1081 3.1959
Llama-2-7B 57.6 ± 4.4 60.0 ± 4.4 58.4 ± 4.4 29.1 ± 4.1 29.9 ± 4.1 36.2 ± 4.3 11.1 ± 1.0 4.6 ± 0.6 1.5 ± 0.3 6.1528 14.8055 3.1198
@@
Stage 1 80.8 ± 3.5 75.2 ± 3.9 65.6 ± 4.3 32.3 ± 4.2 26.8 ± 3.9 26.0 ± 3.9 11.5 ± 1.0 4.2 ± 0.5 2.6 ± 0.6 3.4895 11.4210 3.1341
Stage 2 78.4 ± 3.7 65.6 ± 4.3 74.4 ± 3.9 31.5 ± 4.1 26.0 ± 3.9 28.4 ± 4.0 26.5 ± 1.1 3.4 ± 0.4 15.7 ± 1.0 2.1885 3.8351 1.4055
Stage 2 + parallel 84.0 ± 3.3 66.4 ± 4.2 76.8 ± 3.8 35.4 ± 4.3 27.6 ± 4.0 29.1 ± 4.1 29.1 ± 1.2 4.3 ± 0.5 16.0 ± 1.0 2.1837 3.7615 1.4050
Table 4: Pre-training results for extremely low-resource Finno-Ugric languages. Standard errors are reported for the scores (score ± stderr). For comparison, we report Llama-2 7B and Llammas-base Kuulmets et al. (2024). Stage 1 consists of pre-training with high-resource related languages while Stage 2 additionally includes Võro, Livonian and Komi. Stage 2 + parallel incorporates additional parallel translation data into training.
@@ 5.1 Pre-training
Stage 1 continued pre-training on high-resource supporting languages shows notable improvements in SIB-SMUGRI for Võro and Livonian compared to Llama-2-7B (see Stage 1 in Table 4). There is also a clear reduction in perplexity for Võro and Livonian, while no such improvement is observed for Komi. Similarly, the Llama model fine-tuned for Estonian (Llammas) exhibits lower perplexity than Llama for Livonian and Võro, which are closely related to Estonian. The lack of improvement for Komi may result from its more distant relationship to the other Finno-Ugric languages in the dataset, as well as its use of a different script. These results suggest that related languages generally improve benchmark scores for XLR languages that were not included in the training. For Belebele-SMUGRI and FLORES-SMUGRI there is no visible improvement.
Stage 2 pre-training, which focuses on XLR Finno-Ugric languages, further enhances both perplexity and FLORES-SMUGRI scores (except for Livonian), indicating that the model has acquired generative capabilities for these languages. The performance improvements on the SIB-SMUGRI benchmark are modest for Komi, while Livonian and Võro show a decrease. Belebele scores decline from those of Llama-2-7B, with the exception of Võro, which demonstrates improvement.
Stage 2 + parallel data results in minimal improvements in benchmarks and perplexity. This could indicate that the inclusion of parallel data has a limited impact or that our benchmarks are insufficiently sensitive to capture these effects. Nevertheless, due to the slightly positive influence observed, we will use this setup as the basis for subsequent instruction tuning.
Benchmarks. The current benchmarks may not effectively differentiate between models at this stage, as their small size limits our ability to draw fine-grained conclusions about training strategies. Additionally, the low scores on the Belebele benchmark suggest it may be too challenging for the models. In contrast, the relatively high scores on the SIB-200 benchmark could result from its simplicity, allowing the models to classify texts based on single-word clues rather than a deeper understanding of the language.
@@
GPT-3.5-turbo 45.7 ± 4.4 37.8 ± 4.3 34.6 ± 4.2 81.6 ± 3.5 73.6 ± 4.0 68.8 ± 4.2
GPT-4-turbo 70.1 ± 4.1 40.2 ± 4.3 44.1 ± 4.4 92.0 ± 2.5 72.0 ± 4.0 67.2 ± 4.2
Llammas Kuulmets et al. (2024) 36.2 ± 4.3 32.3 ± 4.2 27.6 ± 4.0 80.8 ± 3.5 78.4 ± 3.7 63.2 ± 4.3
@@
SupInst 42.5 ± 4.4 30.7 ± 4.1 44.1 ± 4.4 86.4 ± 3.1 79.2 ± 3.6 88.8 ± 2.8
SupInst+LLMTrAlpaca 39.4 ± 4.3 35.4 ± 4.3 42.5 ± 4.4 85.6 ± 3.1 81.6 ± 3.5 84.8 ± 3.2
SupInst+TrAlpaca 35.4 ± 4.2 32.3 ± 4.2 40.2 ± 4.3 85.6 ± 3.1 79.2 ± 3.6 85.6 ± 3.1
SupInst+LLMTrAlpaca+TrInst 44.9 ± 4.4 40.9 ± 4.4 44.1 ± 4.4 86.4 ± 3.1 76.0 ± 3.8 78.4 ± 3.7
SupInst+TrAlpaca+TrInst 45.7 ± 4.4 32.3 ± 4.2 44.1 ± 4.4 86.4 ± 3.1 78.4 ± 3.7 78.4 ± 3.7
Table 5: Instruction-tuning evaluation results. Standard errors are reported for the scores (score ± stderr).
@@ 5.2 Instruction-Tuned Models
Examining the scores of commercial systems in Table 5, we observe that these models exhibit at least some understanding of Võro, Livonian, and Komi. Based on benchmark scores, they seem to understand better Võro and Livonian than Komi. This could be explained by the linguistic similarity between these languages and Estonian – an average Estonian speaker can understand most of a Võro text and some of a Livonian text, but not much Komi, as it is more distantly related and written in a different script. The performance of GPT-4-Turbo and GPT-3.5-Turbo aligns with this trend, with scores typically following this order. For instance, GPT-4-Turbo achieves 92% accuracy on the Belebele Estonian benchmark, so it is unsurprising that it also performs well on Võro.
@@
Translation evaluation. When assessing language generation using the FLORES translation benchmark, the results in Table 6 show that GPT-family models can translate from Estonian to Võro quite effectively, suggesting that Võro might have been included in their training data. In contrast, the low BLEU scores for Livonian and Komi indicate very limited translation capabilities. Our LLMs, which were not exposed to translation examples during instruction-tuning, struggle to translate into Võro, Livonian, and Komi. However, they perform better in the reverse direction, even surpassing GPT models for Komi. A closer look reveals that they copy the high-resource language sentences to the output. When the translated Alpaca instructions (TrAlpaca and LLMTrAlpaca) were added, we observed that the models often copied the source text in XLR languages to the output when asked to translate, resulting in lower BLEU scores. This can be addressed by including a small amount of translation data during instruction-tuning (TrInst).
@@ 5.4 Human Evaluation
We select 3 instruction-tuned models for human evaluation: TrAlpaca, LLMTrAlpaca+TrInst and TrAlpaca+TrInst. As a baseline, we use GPT-3.5-turbo, which is freely accessible via a chat interface [^7] https://chatgpt.com/. For each target language, we designed a survey where participants rate the helpfulness of the answer from a randomly selected model on a 5-point Likert scale. Additionally, we ask participants to rate how natural the response sounds in the target language, as Kuulmets et al. (2024) notes that model outputs often sound unnatural in these languages. The surveys were distributed within target language communities via social media and direct outreach to speakers (see Appendix A for the screenshot of the survey). We did not collect any personal data from respondents.
@@ Appendix C Choice of Unimax N
Muennighoff et al. (2023) found that repeating data for 4 epochs is almost as good as new data with improvements continuing beyond 4 epochs. Indeed, we find that with very small datasets, 4 epochs of repetition (Unimax N=4) provides an improvement in perplexity over 1 epoch of data for Stage 2. However, already at 8 epochs, the perplexity drops, suggesting overfitting (see Appendix C). Thus we keep the maximum repetitions at 4 and conclude that the number of repetitions of smaller datasets should be carefully chosen to avoid over- or underfitting.
@@ Appendix E Võro Data Collection
We collect Võro data from Võro language Wikipedia dump Foundation , Corpus of Fiction in Võro and Seto languages [^9] https://metashare.ut.ee/repository/browse/corpus-of-fiction-in-voro-and-seto-languages/2cf454fca0d411eebb4773db
10791bcf485f3f9e7dee447b983f21b074ad3835
@@ Appendix F Instruction-tuning details
The composition of our instruction-tuning dataset is listed in Table 14. Instructions are formatted into a char-format shown in Figure 4. The translation-tuning data format is shown in Figure 5. The fine-tuning loss is calculated on target (assistant) tokens while the rest of the prompt is masked.
@@ Appendix G Instruction translation details
When using our translation-tuned models for translating instructions (for LLMTrAlpaca), the models sometimes leave sentences untranslated in an unpredictable manner. Consequently, we removed examples where the BLEU Papineni et al. (2002) score between the original and translated text exceeds 70. This process may also eliminate some valid examples, as identical texts can occur in both languages.
@@
We find that this concatenation does not harm model translation quality (see Table 16) while improving training effectiveness similar to packing. Additionally, we observed more consistent outputs when translating whole instructions.

### v2

@@ 1 Introduction
We present a case study in developing LLMs for extremely low-resource (XLR) languages, focusing on three Finno-Ugric (smugri [^1] Finno-Ugric translates to Estonian as soome-ugri, to Finnish as suomalais-ugrilaiset, to Võro as soomõ-ugri, and to Livonian as sūomõ-ugrõ, which is why we refer to it as smugri.) languages: Livonian, Võro and Komi. According to Joshi et al. (2020), Livonian is in the lowest category out of 6 (0 / The Left-Behinds) while Võro and Komi are in the second-lowest (1 / The Scraping-Bys). Developing LLMs and other tools for these languages is thus a significant challenge, as well as a vital step for ensuring their digital survival and support.
@@ 3.1 Continued Pre-training
Stage 2 + parallel: making use of parallel translation data. To investigate the role of parallel translation examples in the pre-training data, we incorporate translation examples formatted into various templates, accounting for up to 1% of the Stage 2 character budget (159,712 sentence pairs). We use Unimax with N=1 to balance the budget between language pairs. For further details on the parallel data, we refer the reader to Appendix H. This stage yields the final base model we refer to as Llama-SMUGRI.
@@ 3.2 Instruction-tuning
Supporting instructions. We utilize existing instruction-tuning datasets across multiple languages. For English, Russian, and Finnish, we use Aya Singh et al. (2024), and the highest-rated conversation paths of OASST-2 Köpf et al. (2023). Additionally, we sample 20,000 Estonian instructions from Alpaca-est Kuulmets et al. (2024). Following Kuulmets et al. (2024), we include 5,000 instructions from the FLAN-V2 Longpre et al. (2023) TULU mixture Wang et al. (2023) and 20,000 examples from Alpaca-GPT-4 Peng et al. (2023), to improve cross-lingual knowledge transfer from high-quality English instructions. We refer to this instruction mixture as SupInst (Supporting Instructions). Further details are listed in Appendix F.
@@
To investigate a scenario where a translation model is unavailable, we explore handling translating Alpaca instructions to XLR languages by fine-tuning our base model (Llama-SMUGRI) for the translation task (discussed in §3.3) resulting in LLMTrAlpaca instructions. This is similar to self-translate-train Ri et al. (2024) and self-translate Etxaniz et al. (2024a) concepts with the difference that we add a fine-tuning step to obtain an LLM specialized for translation. Further instruction-translation details and human evaluation of translations are discussed in Appendix G.
@@ 4.2.2 Constructing the Dataset
Instead of human translators, one could use a machine translation system or a proprietary LM for test data generation or translation. We explore both options and find that proprietary models struggle to generate text in our target languages, while even the best machine translation systems produce translations that are often judged inferior to human translations (see Appendix J for further details).
Model SIB-SMUGRI BELEBELE-SMUGRI FLORES-SMUGRI
@@
VRO LIV KPV VRO LIV KPV ET-VRO ET-LIV RU-KPV VRO-ET LIV-ET KPV-RU
Llammas-base 78.4 (3.7) 69.6 (4.1) 64.0 (4.3) 30.7 (4.1) 28.4 (4.0) 32.3 (4.2) 11.5 (0.9) 4.3 (0.5) 1.7 (0.4) 28.7 (1.5) 8.0 (0.8) 2.2 (0.3)
Llama-2-7B 57.6 (4.4) 60.0 (4.4) 58.4 (4.4) 29.1 (4.1) 29.9 (4.1) 36.2 (4.3) 11.1 (1.0) 4.6 (0.6) 1.5 (0.3) 11.3 (0.9) 4.4 (0.6) 2.4 (0.3)
@@
Stage 1 80.8 (3.5) 75.2 (3.9) 65.6 (4.3) 32.3 (4.2) 26.8 (3.9) 26.0 (3.9) 11.5 (1.0) 4.2 (0.5) 2.6 (0.6) 29.6 (1.4) 7.2 (0.7) 4.1 (0.7)
Stage 2 78.4 (3.7) 65.6 (4.3) 74.4 (3.9) 31.5 (4.1) 26.0 (3.9) 28.4 (4.0) 26.5 (1.1) 3.4 (0.4) 15.7 (1.0) 45.3 (1.5) 10.6 (0.9) 18.6 (0.9)
Stage 2 + parallel 84.0 (3.3) 66.4 (4.2) 76.8 (3.8) 35.4 (4.3) 27.6 (4.0) 29.1 (4.1) 29.1 (1.2) 4.3 (0.5) 16.0 (1.0) 48.7 (1.4) 17.6 (1.0) 22.1 (1.3)
Table 4: Pre-training results for extremely low-resource Finno-Ugric languages. Standard errors are reported for the scores in parentheses: score (stderr). For comparison, we report Llama-2 7B and Llammas-base Kuulmets et al. (2024). Stage 1 consists of pre-training with high-resource related languages while Stage 2 additionally includes Võro, Livonian and Komi. Stage 2 + parallel incorporates additional parallel translation data into training.
Model byte-PPL
VRO LIV KPV
Llammas-base 3.3548 12.1081 3.1959
Llama-2-7B 6.1528 14.8055 3.1198
Llama-SMUGRI (ours)
Stage 1 3.4895 11.4210 3.1341
Stage 2 2.1885 3.8351 1.4055
Stage 2 + parallel 2.1837 3.7615 1.4050
Table 5: Pre-training perplexity for extremely low-resource Finno-Ugric languages. For comparison, we report Llama-2 7B and Llammas-base Kuulmets et al. (2024).
@@ 5.1 Pre-training
Stage 1 continued pre-training on high-resource supporting languages shows notable improvements in SIB-SMUGRI for Võro and Livonian compared to Llama-2-7B (see Stage 1 in Table 4). There is also a clear reduction in perplexity (see Table 5) for Võro and Livonian, while no such improvement is observed for Komi. Similarly, the Llama model fine-tuned for Estonian (Llammas-base) exhibits lower perplexity than Llama for Livonian and Võro, which are closely related to Estonian. The lack of improvement for Komi may result from its more distant relationship (see Appendix I) to the other Finno-Ugric languages in the dataset, as well as its use of a different script. These results suggest that related languages generally improve benchmark scores for XLR languages that were not included in the training. For Belebele-SMUGRI, there is no improvement compared to Llama-2-7B, while FLORES-SMUGRI shows improvement only when translating from the XLR languages into the higher-resourced languages.
Stage 2 pre-training, which targets XLR Finno-Ugric languages, further enhances both perplexity and, with the exception of ET-LIV translation, FLORES-SMUGRI scores, indicating that the model has acquired generative capabilities for these languages. The performance improvements on the SIB-SMUGRI benchmark are modest for Komi, while Livonian and Võro show a slight decrease from the previous stage. This stage of training has also failed to improve the Belebele scores.
Stage 2 + parallel data results in minimal improvements in benchmarks and perplexity, with the exception of translation tasks from the XLR languages showing larger gains. This indicates that the inclusion of parallel data has a limited impact or that our benchmarks are insufficiently sensitive to capture these effects. Nevertheless, due to the slightly positive influence observed, we will use this setup as the basis for subsequent instruction tuning.
Benchmarks. The current benchmarks may not effectively differentiate between models at this stage, as their small size and high standard errors limit our ability to draw fine-grained conclusions about training strategies. Additionally, the low scores on the Belebele benchmark suggest it may be too challenging for the models. In contrast, the relatively high scores on the SIB-200 benchmark could result from its simplicity, allowing the models to classify texts based on single-word clues rather than a deeper understanding of the language. Designing automatic benchmarks with an appropriate difficulty level and relevant context for these languages is an important challenge for future research.
@@
GPT-3.5-turbo 45.7 (4.4) 37.8 (4.3) 34.6 (4.2) 81.6 (3.5) 73.6 (4.0) 68.8 (4.2)
GPT-4-turbo 70.1 (4.1) 40.2 (4.3) 44.1 (4.4) 92.0 (2.5) 72.0 (4.0) 67.2 (4.2)
Llammas Kuulmets et al. (2024) 36.2 (4.3) 32.3 (4.2) 27.6 (4.0) 80.8 (3.5) 78.4 (3.7) 63.2 (4.3)
@@
SupInst 42.5 (4.4) 30.7 (4.1) 44.1 (4.4) 86.4 (3.1) 79.2 (3.6) 88.8 (2.8)
SupInst+LLMTrAlpaca 39.4 (4.3) 35.4 (4.3) 42.5 (4.4) 85.6 (3.1) 81.6 (3.5) 84.8 (3.2)
SupInst+TrAlpaca 35.4 (4.2) 32.3 (4.2) 40.2 (4.3) 85.6 (3.1) 79.2 (3.6) 85.6 (3.1)
SupInst+LLMTrAlpaca+TrInst 44.9 (4.4) 40.9 (4.4) 44.1 (4.4) 86.4 (3.1) 76.0 (3.8) 78.4 (3.7)
SupInst+TrAlpaca+TrInst 45.7 (4.4) 32.3 (4.2) 44.1 (4.4) 86.4 (3.1) 78.4 (3.7) 78.4 (3.7)
Table 6: Instruction-tuning evaluation results. Standard errors are reported for the scores: score (stderr).
@@ 5.2 Instruction-Tuned Models
Examining the scores of commercial systems in Table 6, we observe that these models exhibit at least some understanding of Võro, Livonian, and Komi. Based on benchmark scores, they seem to understand Võro and Livonian better than Komi. This could be explained by the linguistic similarity between these languages and Estonian – an average Estonian speaker can understand most of a Võro text and some of a Livonian text, but not much Komi, as it is more distantly related and written in a different script. The performance of GPT-4-turbo and GPT-3.5-turbo aligns with this trend, with scores typically following this order. For instance, GPT-4-turbo achieves 92% accuracy on the Belebele Estonian benchmark, so it is unsurprising that it also performs well on Võro.
@@
Translation evaluation. When assessing language generation using the FLORES translation benchmark, the results in Table 7 show that GPT-family models can translate from Estonian to Võro quite effectively, suggesting that Võro might have been included in their training data. In contrast, the low BLEU scores for Livonian and Komi indicate very limited translation capabilities. Our LLMs, which were not exposed to translation examples during instruction-tuning, struggle to translate into Võro, Livonian, and Komi. However, they perform better in the reverse direction, even surpassing GPT models for Komi. A closer look reveals that they copy the high-resource language sentences to the output when translating to the low-resource languages. When the translated Alpaca instructions (TrAlpaca and LLMTrAlpaca) were added, we observed that when asked to translate from the low-resource languages, the models often copied the source text to the output as well, resulting in lower BLEU scores. This can be addressed by including a small amount of translation data during instruction-tuning (TrInst).
@@ 5.4 Human Evaluation
We select 3 instruction-tuned models for human evaluation: TrAlpaca, LLMTrAlpaca+TrInst and TrAlpaca+TrInst. As a baseline, we use GPT-3.5-turbo, which was freely accessible via a chat interface [^7] https://chatgpt.com/ at the time of the evaluation. For each target language, we designed a survey where participants rate the helpfulness of the answer from a randomly selected model on a 5-point Likert scale. Additionally, we ask participants to rate how natural the response sounds in the target language, as Kuulmets et al. (2024) notes that model outputs often sound unnatural in these languages. The surveys were distributed within target language communities via social media and direct outreach to speakers (see Appendix A for the screenshot of the survey). We did not collect any personal data from respondents.
@@ 6 Conclusion
A limitation of our instruction-tuning process is that we only used machine-translated instructions for the XLR languages. As a result, some of these instructions were of low quality, potentially affecting the overall performance and reliability of the fine-tuned models.
@@ Appendix C Choice of Unimax N
Muennighoff et al. (2023) found that repeating data for 4 epochs is almost as good as new data with improvements continuing beyond 4 epochs for pre-training LLMs. We find that for continued pre-training with very small datasets and small data budgets, 4 epochs of repetition (Unimax N=4) provides an improvement in perplexity over 1 epoch of data for Stage 2. However, already at 8 epochs, the perplexity drops, suggesting overfitting (see Appendix C). Thus we keep the maximum repetitions at 4 and conclude that the number of repetitions of smaller datasets should be carefully chosen to avoid over- or underfitting.
@@ Appendix E Võro Data Collection
We collect Võro data from Võro language Wikipedia dump Foundation , Corpus of Fiction in Võro and Seto languages [^9] https://metashare.ut.ee/repository/browse/corpus-of-fiction-in-voro-and-seto-languages/2cf454fca0d411eebb47
73db10791bcf485f3f9e7dee447b983f21b074ad3835
@@ Appendix F Instruction-tuning details
The composition of our instruction-tuning dataset is listed in Table 15. Instructions are formatted into a chat-format shown in Figure 5. The translation-tuning data format is shown in Figure 6. The fine-tuning loss is calculated on target (assistant) tokens while the rest of the prompt is masked.
@@ Appendix G Instruction translation details
When using our translation-tuned models for translating instructions (for LLMTrAlpaca), the models sometimes leave sentences untranslated in an unpredictable manner. Consequently, we removed examples where the BLEU Papineni et al. (2002) score between the original and translated text exceeds 70. This process may also eliminate some valid examples, as identical texts can occur in some cases.
@@
We find that this concatenation does not harm model translation quality (see Table 17). Additionally, we observed more consistent outputs when translating whole instructions.
To get a glimpse into the quality of the translations, we conducted a small-scale human evaluation with native speakers or, in the case of Livonian a fluent speaker, due to the lack of native speakers. Given the original and the translation of 20 randomly chosen instructions, which stay the same across translation models, the evaluators were asked to rate the Fluency (How fluent and natural does the translation sound in the target language?) and Consistency (Does the translation preserve the meaning and intent of the source text?) on a 5-point Likert scale (see Table 18 for evaluation guidelines). We also asked them to report if the resulting instruction-input-output triplet does not form a correct instruction, e.g. the output does not satisfy the instruction.
We report the evaluation summary in Table 19 for Neurotõlge and Llama-SMUGRI-translate. In general, Llama-SMUGRI-translate produces better results for Võro and Livonian, achieving higher fluency and consistency scores than Neurotõlge on average. For an acceptable instruction translation, we would like the fluency and consistency to be at least 3, and the resulting instruction should still be correct. We see that for Neurotõlge, this is only achieved for 45% of evaluated instructions in Võro and Livonian, while for Komi 80% of instructions are acceptable. For Llama-SMUGRI-translate, we see this condition is satisfied for 70%, 80%, and 65% of translation for Komi, Võro, and Livonian, respectively. While this shows that machine translation can be a feasible option in some cases, it is far from ideal for instructions. We also report histograms of human ratings in Figure 4.
Fluency
5 The translation is perfectly fluent, with no grammatical errors, unnatural phrasing, or awkward expressions.
4 The translation is mostly fluent, with minor grammatical or stylistic issues that do not affect readability.
3 The translation is somewhat fluent, but noticeable issues (e.g., awkward phrasing or grammatical errors) hinder smooth reading.
2 The fluency of the translation is poor, with significant issues that make it difficult to understand in parts.
1 The translation is completely unnatural or ungrammatical, making it incomprehensible.
Consistency
5 The translation fully preserves the meaning of the source text, with no omissions, additions, or distortions.
4 The translation preserves the overall meaning, but there are minor inaccuracies or nuances lost.
3 The translation conveys the general meaning, but there are noticeable issues (e.g., omissions or slight distortions).
2 The translation distorts the meaning significantly or omits important details, making it partially inaccurate.
1 The translation fails to convey the meaning of the source text entirely.
Table 18: Instruction translation human evaluation guidelines.
[…]
Neurotõlge
KPV 3.55 3.5 5% 80%
VRO 2.75 3.2 0% 45%
LIV 2.8 2.85 15% 45%
Llama-SMUGRI-translate
KPV 3.35 3.35 5% 70%
VRO 3.5 3.9 0% 80%
LIV 3.2 3.2 20% 65%
[…]
Figure 4: Neurotõlge and Llama-SMUGRI-translate instruction translation human evaluation scores.